Kurt's dog eats 3 scoops of dog food each day. If you write down the total number of scoops of food the dog has consumed over time, what kind of sequence will you have?

Arithmetic geometric both neither

Answer by jim_thompson5910(35256)

Day 1 = dog eats 3 scoops so far
Day 2 = dog eats 6 scoops so far (3+3 = 6)
Day 3 = dog eats 9 scoops so far (6+3 = 9)
Day 4 = dog eats 12 scoops so far (9+3 = 12)
The pattern keeps going

The sequence {3, 6, 9, 12, ...} is an arithmetic sequence where the first term is a1 = 3 and the common difference is 3.